### v1

# MFCLIP: Multi-modal Fine-grained CLIP for Generalizable Diffusion Face Forgery Detection

Abstract
The rapid development of photo-realistic face generation methods has raised significant concerns in society and academia, highlighting the urgent need for robust and generalizable face forgery detection (FFD) techniques. Although existing approaches mainly capture face forgery patterns using image modality, other modalities like fine-grained noises and texts are not fully explored, which limits the generalization capability of the model. In addition, most FFD methods tend to identify facial images generated by GAN, but struggle to detect unseen diffusion-synthesized ones. To address the limitations, we aim to leverage the cutting-edge foundation model, contrastive language-image pre-training (CLIP), to achieve generalizable diffusion face forgery detection (DFFD). In this paper, we propose a novel multi-modal fine-grained CLIP (MFCLIP) model, which mines comprehensive and fine-grained forgery traces across image-noise modalities via language-guided face forgery representation learning, to facilitate the advancement of DFFD. Specifically, we devise a fine-grained language encoder (FLE) that extracts fine global language features from hierarchical text prompts. We design a multi-modal vision encoder (MVE) to capture global image forgery embeddings as well as fine-grained noise forgery patterns extracted from the richest patch, and integrate them to mine general visual forgery traces. Moreover, we build an innovative plug-and-play sample pair attention (SPA) method to emphasize relevant negative pairs and suppress irrelevant ones, allowing cross-modality sample pairs to conduct more flexible alignment. Extensive experiments and visualizations show that our model outperforms the state of the arts on different settings like cross-generator, cross-forgery, and cross-dataset evaluations.
Index Terms:
Diffusion face forgery detection, Transformer, CLIP, Image-noise fusion, Sample pair attention.
I Introduction
Photorealistic generation models like generative adversarial networks (GANs) [1, 2, 3] and denoising diffusion probabilistic models (DDPM) [4], have reached unprecedented progress in synthesizing highly realistic facial images. Thus, advancing face forgery detection (FFD) becomes a critical and urgent demand.
Figure 1: (a) The visualization of the richest or poorest patch high-frequency noises produced by steganalysis rich model (SRM) [5]. The first row represents the RGB images. The second and third rows display the richest or poorest patches and corresponding noises of the real and fake samples of various manipulations, respectively. Inspired by [6], we split an image into non-overlapping patches, and select the richest patch as well as the poorest patch, respectively. The richest patch is defined as the region with the largest texture diversity, where the texture diversity is measured by the homogeneity of the gray-level co-occurrence matrix (GLCM). (b) Comparison of cross-modal feature alignment between CLIP and MFCLIP.
Current FFD methods mainly involve three categories, spatial-based models [7, 8, 9, 10, 11], frequency-based models [12, 13, 14], and vision-language-based models [15, 16]. Spatial-based methods [7, 10] tend to mine domain-specific features in GAN-synthesized images, such as identity and background, which leads to poor generalization to unseen facial images generated by diffusion models. Frequency-based methods [17, 12, 18] aim to explore common face forgery traces in the frequency domain, since prior investigations [19, 20] have demonstrated that forgery artifacts produced by various synthesis approaches are mainly concentrated in the high-frequency domain, but they are inclined to capture high-frequency information at a coarse-grained level. As Fig. 1 (a) shows, face images synthesized by advanced diffusion-based models are so realistic that high-frequency noises are rarely observed. Only coarse-grained extraction of high-frequency noises is insufficient to dig comprehensive and universal forgery artifacts. Vision-language-based methods [15, 16] intend to study general face manipulation patterns using text-guided image forgery representation learning, but they struggle to introduce hierarchical fine-grained text prompts and achieve flexible cross-modal alignment, to facilitate the generalization to face images synthesized using diffusion models.
Figure 2: An overview of the proposed MFCLIP. (a) The vanilla CLIP model tends to extract coarse-grained text embeddings, and only focuses on image-text modality, which leads to poor generalization to diffusion face images. (b) Our MFCLIP model captures fine-grained semantic information, and mines general multi-modal manipulated patterns via text-guided forgery representation learning, to achieve strong generalization to diffusion face images.
Based on the aforementioned discussion, we propose to introduce fine-grained text prompts to boost the learning of common visual forgery patterns across noise and image modalities via adaptive cross-modal feature alignment, to achieve generalizable diffusion face forgery detection (DFFD). In this paper, inspired by the cutting-edge contrastive language-image pre-training (CLIP) [21] model, we design a multi-modal fine-grained CLIP (MFCLIP). Unlike CLIP which is pre-trained on large-scale, general natural images, and then fine-tuned to improve inference on downstream tasks like FFD, our MFCLIP model is trained from scratch in an end-to-end manner using facial images generated by various generators like diffusion, to facilitate the advancement of DFFD. Specifically, our MFCLIP model primarily differs from CLIP in the following aspects (see Fig. 2): First, we observe that there are significant differences between authentic and forgery facial SRM [5] noises from the richest patches (see Fig. 1), compared to the poorest patches. Specifically, SRM noises extracted from the richest patch are visually evident in real images, but not obvious in fake ones. Besides, the SRM noise in the GAN-generated richest patch is more noticeable than that in the diffusion-generated one. Therefore, we design a multi-modal vision encoder (MVE) with a noise encoder, to study the fine-grained and discriminative noise forgery patterns extracted from the richest patches. Second, we devise a fine-grained text generator (FTG) to build the hierarchical text prompts, and a fine-grained language encoder (FLE) to capture detailed global relations among text prompts. Finally, inspired by the CLIP, we aim to enhance visual forgery representations through feature alignment between cross-modal sample pairs. However, we notice that CLIP tends to maximize the distance between relevant negative pairs (see Fig. 1 (b)), which is regarded as a sub-optimal feature alignment solution since the relevant negative pairs should be closer to each other in the feature space. To address the limitation, we design a plug-and-play sample pair attention (SPA) module to flexibly emphasize relevant negative pairs and suppress irrelevant ones. To sum up, the contributions of our work are as follows:
$\bullet$ We propose a novel MFCLIP model for generalizable DFFD, which incorporates fine-grained noises extracted from the richest patches with global image forgery artifacts, as well as enhances visual features across image-noise modalities via text-guided face forgery representation learning.
$\bullet$ We devise an innovative plug-and-play sample pair attention (SPA) method to adaptively emphasize relevant negative pairs and suppress irrelevant ones, which can be integrated into any vision-language-based models with only a slight growth in computational costs.
$\bullet$ Extensive experiments and visualizations show that our method outperforms the state of the arts on various protocols such as cross-generator evaluation, cross-forgery evaluation, and cross-dataset evaluation.
The remainder of the paper is organized as follows: Section II introduces related work, and Section III explains the proposed MFCLIP method. Section IV describes the experiments, including experimental settings and comparison with the state of the art. Section V analyses the results of the ablation study. Section VI presents the visualizations, and the conclusion and limitations are discussed in Section VII.
II Related Work
Our research mainly involves two aspects: face forgery detection and vision-language models. In the following subsections, we discussed the two points, separately.
II-A Face Forgery Detection
Existing efforts have achieved considerable progress in the field of FFD. Some methods focus on forgery artifacts in the spatial domain. Rossler et al.[22] employed an Xception pre-trained on ImageNet, to capture local spatial forgery artifacts. To explore comprehensive relations among image patches, a convolutional vision transformer (CViT) model [10] is designed to combine CNN with vision transformer (ViT) [8] for FFD, to mine global forgery traces in the spatial domain. Diffusion reconstruction error (DIRE) [9] utilizes the discrepancy between an input image and corresponding reconstruction for diffusion-generated image detection. In addition to employing spatial forgery patterns for FFD, there are also some frequency-aware approaches [12, 13, 14]. SFDG [14] is proposed to mine the relation-aware representations in spatial and frequency domains using dynamic graph learning. FreqNet [12] is devised to focus on high-frequency forgery traces across spatial and channel dimensions. To capture diverse forgery patterns across various modalities, TwoStream [13] is designed to integrate the high-frequency noise features with RGB content for generalized FFD. Recently, VLFFD [15] is proposed to acquire more generalization and interpretability for FFD, where fine-grained text prompts and coarse-grained original data are collaboratively employed to guide the coarse-and-fine co-training network learning. DD-VQA [16] introduces a multi-modal transformer framework, which boosts the learning of face forgery representations using a text and image contrastive learning scheme, to facilitate the FFD. By contrast, since there are significant discrepancies between authentic and forgery facial noise images extracted by SRM from the richest patches, we solely extract the SRM noises from the single richest patch, and combine them with global spatial forgery patterns to learn more general embeddings for DFFD. Furthermore, we boost common forged representations across image and noise modalities via fine-grained vision-language contrastive learning.
II-B Vision-Language Models
Different from conventional image-based models that consist of an image feature encoder and a classifier to predict a fixed set of predefined sample categories, the vision-language models [23, 24, 25] such as contrastive language-image pre-training (CLIP) [21] simultaneously train an image encoder and a text encoder, to match image and text pairs from training datasets. CLIP facilitates inference on downstream tasks using a zero-shot linear classifier embedded with the class names or descriptions from the target dataset, showing robust image representations across various domains, including object detection and image generation. To circumvent the labour-intensive and inconsistent performance associated with prompt engineering for text encoders, CoOp [26] models the context words of a prompt using learnable vectors, and places the class token (i.e., names or descriptions) at the end or middle position. CFPL [27] is designed to study the various semantic prompts conditioned on different visual features for generalizable face anti-spoofing. Unlike traditional CLIP-based methods, our model is capable of capturing visual face forgery embeddings across image-noise modalities, and aligning the cross-modality features using the SPA method, adaptively.
Figure 3: The framework of MFCLIP. We first generate the hierarchical text prompts through the FTG module, to build the image-text pairs. We then send them to MVE to explore general visual forgery representations across image-noise modalities, and FLE to capture the detailed text embeddings, respectively. Subsequently, they are transferred to a SPA module to align features between cross-modal sample pairs, adaptively. Meanwhile, the visual features are fed into the predictor to predict text embeddings, and the MLP head to acquire the final prediction, respectively.
III Methodology
III-A Method Overview
To capture general multi-modal visual forgery patterns for generalizable DFFD, we design the multi-modal fine-grained CLIP model, to conduct the fine-grained text-guided visual forgery representation learning. As Fig. 3 shows, MFCLIP mainly consists of four key modules: fine-grained text generator (FTG), multi-modal vision encoder (MVE), fine-grained language encoder (FLE), and the sample pair attention (SPA) module. During the training phase, given a batch of input facial images $X$, MFCLIP generates the corresponding text prompts $T$ of each image from the GenFace dataset $D$ via the FTG module, and obtains the image-text pairs $(X,T)\in D$ with one-hot labels $y\in\big{\{}[0,1]^{T},[1,0]^{T}\}$. MFCLIP then captures the multi-modal visual forgery embeddings $X_{\text{v}}$ through the MVE module, and extracts abundant fine-grained language embeddings $T_{\text{l}}$ via the FLE module, respectively. Thereafter, $X_{\text{v}}$ and $T_{\text{l}}$ are transmitted to the SPA module to adaptively emphasize and suppress sample pairs, to generate $X_{\text{spa}}$. Meanwhile, $X_{\text{v}}$ is fed into a predictor to predict language features $T_{\text{l}}^{\text{pre}}$. Finally, the multilayer perceptron (MLP) head consisting of a full connection layer generates the final predictions $y_{\text{pre}}$. The kullback-leibler (KL) divergence loss function $\mathcal{L}_{\text{kl}}$ is leveraged to measure the difference between predicted language features $T_{\text{l}}^{\text{pre}}$ and the true language embeddings $T_{\text{l}}$. The cross-modality contrastive (CMC) loss $\mathcal{L}_{\text{cmc}}$ is used to align features between vision-language sample pairs. The cross-entropy (CE) loss $\mathcal{L}_{\text{ce}}$ is utilized to compute the discrepancy between the predicted label $y_{\text{pre}}$ and the ground truth label $y$. That is,
$\text{MFCLIP}(X,T)$ $=\text{SPA}(\text{MVE}(X),\text{FLE}(T))$
$=\text{SPA}(X_{\text{v}},T_{\text{l}})=X_{\text{spa}}$ (1)
$\mathcal{L}(X,T)$ $=\ \mathcal{L}_{\text{kl}}\ (T_{\text{l}}^{\text{pre}},T_{\text{l}})+\mathcal{L}_{\text{cmc}}(X_{\text{spa}})+\mathcal{L}_{\text{ce}}(y_{\text{pre}},y)$ (2)
During the inference phase, MFCLIP adopts MVE to generate the visual forgery pattern across image-noise modalities, which is then fed into the MLP head to yield the final prediction. It can be expressed formally as follows:
$\text{MFCLIP}(X)$ $=\text{MLP}(\text{MVE}(X))$
$=\text{MLP}(X_{\text{v}})$
$=y_{\text{pre}}$ (3)
III-B Fine-grained Text Generator
To create the hierarchical fine-grained text prompts, we design the fine-grained text generator (FTG) module. As Fig. 4 shows, we generate corresponding text prompts for each image based on the hierarchical fine-grained labels provided by the GenFace dataset, to build the image-text pairs. Specifically, we create corresponding text descriptions for each hierarchical level of an image. At level 1, we formulate the texts “a photo of a real face” or “a photo of a fake face”. The second level describes the forged images into three types, i.e., “a photo of an identity swapped face”, “a photo of an attributed manipulated face”, and “a photo of an entire synthesized face”. We then generate the text prompts “a photo generated by the diffusion-based model” or “a photo generated by the GAN-based model” at level 3. The level 4 refers to the description of the specific generators. We then feed image-text pairs to MVE and FLE, respectively.
Figure 4: The schematic diagram of the fine-grained text generator module.
III-C Multi-modal Vision Encoder
Unlike conventional multi-modal FFD methods [28, 17] that mainly concentrate on the interaction between the frequency or noise and RGB information at a coarse-grained level, we design the multi-modal vision encoder (MVE) to capture global spatial forgery traces as well as fine-grained noise manipulated patterns extracted from the richest patches, and integrate them in a simple and effective way, to achieve generalizable DFFD.
As illustrated in Fig. 3, MVE mainly consists of the image encoder (IE) and noise encoder (NE). IE is composed of a convolutional vision transformer (CViT) [10] model, and NE contains a CNN backbone with stacked convolutional layers and a noise transformer encoder (NoT) with $B$ transformer blocks, $\text{TB}_{j}^{\text{n}}$, $j=1,2,…,B$.
[…]
$N_{1}^{\text{tra}}=N_{\text{tok}}+P_{\text{n}}.$ (4)
Thereafter, it is sequentially fed into $B$ blocks, i.e.,
$\text{NoT}(N_{1}^{\text{tra}})$ $=\text{TB}_{B}^{\text{n}}\circ\text{TB}_{B-1}^{\text{n}}\circ\cdots\circ\text{TB}_{2}^{\text{n}}\circ\text{TB}_{1}^{\text{n}}(N_{1}^{\text{tra}})$
$=\text{TB}_{B}^{\text{n}}\circ\text{TB}_{B-1}^{\text{n}}\circ\cdots\circ\text{TB}_{2}^{\text{n}}(N_{2}^{\text{tra}})$
$=\cdots=\text{TB}_{B}^{\text{n}}\circ\text{TB}_{B-1}^{\text{n}}(N_{B-1}^{\text{tra}})$
$=\text{TB}_{B}^{\text{n}}(N_{B}^{\text{tra}})$
$=N_{\text{NoT}},$ (5)
where $\circ$ denotes function decomposition. NE captures the extensive noise forgery traces $N_{\text{n}}\in\mathbb{R}^{b\times d}$ using the class token in $N_{\text{NoT}}$. MVE finally yields the visual forgery features $X_{\text{v}}$ across image-noise modalities by adding $X_{\text{i}}$ with $N_{\text{n}}$ element-wisely, i.e., $X_{\text{v}}=X_{\text{i}}+N_{\text{n}}$, which are then fed into the SPA module, predictor, and the MLP head, respectively.
Figure 5: The pipeline of the patch selector module. The smaller the homogeneity score, the richer the texture diversity.
III-D Fine-grained Language Encoder
To extract hierarchical fine-grained text embeddings, we devise the fine-grained language encoder (FLE), which consists of a text encoder (TE) with $L$ transformer blocks ${\text{TB}_{j}^{\text{t}}}$, $j=1,2,…,L$. Specifically, given a batch of text prompts $T$, each of which $T_{m}$ contains four sentences $\{T_{mi}^{\text{sen}}\}_{i=1}^{4}$. For each sentence $T_{mi}^{\text{sen}}$, we use the tokenizer [21] to obtain a sequence of word tokens $T_{mi}^{\text{tok}}\in\mathbb{R}^{77}$ , which is then mapped to the word embedding vector $T_{mi}^{\text{emb}}\in\mathbb{R}^{77\times d}$ to obtain ${\ T}_{i}^{\text{emb}}\in\mathbb{R}^{b\times 77\times d}$. We concatenate four sentences to yield ${T}^{\text{emb}}\in\mathbb{R}^{b\times 308\times d}$, which is then added with position embedding $P_{\text{t}}\in\mathbb{R}^{b\times 308\times d}$, i.e., $T_{1}^{\text{tra}}={\ T}^{\text{emb}}+P_{\text{t}}$. Thereafter, it is sequentially fed into $L$ blocks, i.e.,
$\text{TE}(T_{1}^{\text{tra}})$ $=\text{TB}_{L}^{\text{t}}\circ\text{TB}_{L-1}^{\text{t}}\circ\cdots\circ\text{TB}_{2}^{\text{t}}\circ\text{TB}_{1}^{\text{t}}(T_{1}^{\text{tra}})$
$=\text{TB}_{L}^{\text{t}}\circ\text{TB}_{L-1}^{\text{t}}\circ\cdots\circ\text{TB}_{2}^{\text{t}}(T_{2}^{\text{tra}})$
$=\cdots=\text{TB}_{L}^{\text{t}}\circ\text{TB}_{L-1}^{\text{t}}(T_{L-1}^{\text{tra}})$
$=\text{TB}_{L}^{\text{t}}(T_{L}^{\text{tra}})$
$=T_{\text{fg}}.$ (6)
FLE outputs the fine-grained global language embeddings $T_{\text{l}}\in\mathbb{R}^{b\times d}$ using the last class embedding in $T_{\text{fg}}$, which are then transmitted to the SPA module.
III-E Sample Pair Attention
To enhance cross-modal feature alignment flexibly, we propose the sample pair attention (SPA) module. Unlike vanilla CLIP [21] which adjusts features using the static alignment mechanism, our MFCLIP dynamically emphasizes relevant sample pairs and suppresses irrelevant ones, enabling more flexible and effective cross-modal representation learning in a simple and efficient way.
As Fig. 4 illustrates, given a batch of visual forgery features $X_{\text{v}}\in\mathbb{R}^{b\times d}$ and corresponding fine-grained language representations $T_{\text{l}}\in\mathbb{R}^{b\times d}$, we obtain a batch of $b$ vision-language sample pairs $\left\{{(X}_{\text{v}}^{i}{,T}_{\text{l}}^{i})\in\mathbb{R}^{d}\right\}_{i=1}^{b}$. CLIP uses the InfoNCE loss to conduct the cross-modal feature alignment, which pulls the positive pairs together while pushing negative ones apart. However, if some correlated negative pairs are pushed away, noises may be introduced. To address this limitation, MFCLIP generates the vision-language cosine similarity matrix via the SPA method, to emphasize or suppress sample pairs, adaptively.
$S_{\text{v2l}}(X_{\text{v}},T_{\text{l}})={(X}_{\text{v}}T_{\text{l}}^{T})\odot\sigma(A)\in\mathbb{R}^{d\times d},$ (7)
where $A\ \in\mathbb{R}^{d\times d}$ is a learnable weight matrix, $\odot$ denotes a element-wise product, and $\sigma$ is a sigmoid function. Similarly, MFCLIP generates the language-vision cosine similarity matrix,
$S_{\text{l2v}}({T_{\text{l}},X}_{\text{v}})={(T}_{\text{l}}X_{\text{v}}^{T})\odot\sigma(A)\in\mathbb{R}^{d\times d}.$ (8)
That is, for the $i$-th pair, the normalized vision-language similarity vector and the language-vision one are defined as below:
$S_{\text{v2l}}^{ij}(X_{\text{v}},T_{\text{l}})=\frac{\text{exp}(\text{sim}{(X}_{\text{v}}^{i}{,T}_{\text{l}}^{j})/\tau)}{\sum_{j=1}^{b}{\text{exp}(\text{sim}{(X}_{\text{v}}^{i}{,T}_{\text{l}}^{j})/\tau)}}\odot\sigma(A^{ij}),$ (9)
$S_{\text{l2v}}^{ij}({T_{\text{l}},X}_{\text{v}})=\frac{\text{exp}(\text{sim}{(T}_{\text{l}}^{i}{,X}_{\text{v}}^{j})/\tau)}{\sum_{j=1}^{b}{\text{exp}(\text{sim}{(T}_{\text{l}}^{i}{,X}_{\text{v}}^{j})/\tau)}}\odot\sigma(A^{ij}),$ (10)
where $\tau$ is a learnable temperature parameter initialized with 0.07, and the function sim(·) performs a dot product to compute the similarity scores.
III-F Loss Function
Cross-modal contrastive loss. To minimize the distance between positive pairs, while pushing negative pairs away, we introduce the cross-modal contrastive loss function for feature alignment. The one-hot label $y_{\text{pa}}$ of the $i$-th pair is denoted as $y_{\text{pa}}^{i}=\{{y_{\text{pa}}^{ij}}\}_{j=1}^{b}$, $y_{\text{pa}}^{ii}=1$, $y_{\text{pa}}^{ij,i\neq j}=0$,
$\mathcal{L}_{\text{v2l}}\left(X,T\right)$ $=\frac{1}{b}\sum_{i=1}^{b}\text{CE}(S_{\text{v2l}}^{i}\left(X_{\text{v}},T_{\text{l}}\right),y_{\text{pa}}^{i})$
$=\frac{1}{b}\sum_{i=1}^{b}{-{y_{\text{pa}}^{i}}^{T}\text{log}(}S_{\text{v2l}}^{i}\left(X_{\text{v}},T_{\text{l}}\right)),$ (11)
$\mathcal{L}_{\text{l2v}}(T,X)$ $=\frac{1}{b}\sum_{i=1}^{b}{\text{CE}(S_{\text{l2v}}^{i}({T_{\text{l}},X}_{\text{v}}),y_{\text{pa}}^{i})}$
$=\frac{1}{b}\sum_{i=1}^{b}{-{y_{\text{pa}}^{i}}^{T}\text{log}(}S_{\text{l2v}}^{i}({T_{\text{l}},X}_{\text{v}})),$ (12)
where CE is cross-entropy. The final cross-modal contrastive loss $\mathcal{L}_{\text{cmc}}=\ (\mathcal{L}_{\text{v2l}}+\mathcal{L}_{\text{l2v}})/2$.
Kullback-leibler divergence loss. In order to narrow the gap between the predicted language distribution and the true language distribution, we introduce the kullback-leibler divergence loss. Specifically, we send visual forgery embeddings $X_{\text{v}}$ to the predictor consisting of a full connection layer, to produce the predicted language representations $T_{\text{l}}^{\text{pre}}\in\mathbb{R}^{b\times d}$. We leverage the Kullback-leibler divergence loss to bring the predicted language features closer to the authentic ones,
$\mathcal{L}_{\text{kl}}(X,T)=\frac{1}{b}\sum_{i=1}^{b}{\delta{(T_{\text{l}}^{i})}^{T}\text{log}\frac{{\delta(T}_{\text{l}}^{i})}{\delta(T_{\text{l}}^{\text{pre}i})}},$ (13)
where $\delta$ is the softmax function with temperature 0.5, to smooth representations.
Cross-entropy loss. To minimize the distance between the predicted label and the true label, we use the cross-entropy loss function. In detail, we transfer $X_{\text{v}}$ to the MLP head with a full connection layer, to generate the final prediction $y_{\text{pre}}$. We use the cross-entropy loss as follows:
$\mathcal{L}_{\text{ce}}(X)=\ \frac{1}{b}\sum_{i=1}^{b}{-{y^{i}}^{T}\text{log}(}y_{\text{pre}}^{i}).$ (14)
Figure 6: Comparison of the performance of detectors on various generators. We averaged the AUC scores of the model listed in Table III on the images from the same generator.
TABLE I: Cross-forgery generalization. ACC and AUC scores (%) on remaining manipulations, after training using one manipulation.
Training Set Model Testing Set
EFS AM FS
ACC AUC ACC AUC ACC AUC
EFS Xception [7] - - 50.00 63.14 68.06 79.52
ViT [8] - - 54.69 65.86 53.13 61.43
CViT [10] - - 50.02 63.53 72.79 73.82
DIRE [9] - - 50.03 76.14 74.03 77.89
FreqNet [12] - - 50.00 75.41 76.48 69.62
CLIP [21] - - 56.05 64.32 53.06 61.40
FatFormer [29] - - 55.89 66.90 56.21 64.89
VLFFD [15] - - 55.31 73.89 69.54 70.31
DD-VQA [16] - - 55.96 74.02 70.23 80.02
MFCLIP (Ours) - - 58.05 78.76 76.88 81.99
AM Xception [7] 50.20 51.45 - - 50.11 54.57
ViT [8] 50.29 60.37 - - 50.19 55.04
CViT [10] 50.15 70.32 - - 50.02 60.74
DIRE [9] 51.14 72.41 - - 51.24 70.45
FreqNet [12] 50.88 74.68 - - 50.33 76.34
CLIP [21] 52.48 60.96 - - 51.01 55.21
FatFormer [29] 51.67 63.90 - - 51.82 59.03
VLFFD [15] 52.06 68.98 - - 51.52 74.40
DD-VQA [16] 52.15 73.10 - - 52.07 77.50
MFCLIP (Ours) 53.29 87.76 - - 52.61 80.31
FS Xception [7] 50.42 76.48 53.75 75.62 - -
ViT [8] 51.09 69.16 52.37 78.11 - -
CViT [10] 50.22 73.88 49.98 73.75 - -
DIRE [9] 54.06 79.65 52.13 78.32 - -
FreqNet [12] 53.46 73.68 51.97 74.18 - -
CLIP [21] 52.10 71.47 50.16 62.27 - -
FatFormer [29] 54.06 74.78 53.81 65.89 - -
VLFFD [15] 55.38 80.74 54.60 77.14 - -
DD-VQA [16] 56.02 81.10 54.79 76.05 - -
MFCLIP (Ours) 60.08 84.93 62.58 80.38 - -
TABLE II: Cross-generator evaluation on FS. ACC and AUC scores (%) on remaining generators, after training using one generator.
Training Set Model Testing Set
DiffFace FSLSD FaceSwapper
ACC AUC ACC AUC ACC AUC
DiffFace Xception [7] - - 50.00 48.45 50.00 83.60
ViT [8] - - 52.27 54.47 65.58 86.02
CViT [10] - - 50.00 49.28 50.04 79.17
DIRE [9] - - 50.00 55.49 50.00 88.01
FreqNet [12] - - 49.75 44.42 49.65 69.93
CLIP [21] - - 51.06 55.37 70.54 92.62
FatFormer [29] - - 51.95 57.35 71.87 93.76
VLFFD[15] - - 52.44 56.03 72.97 94.03
DD-VQA [16] - - 52.51 57.89 73.74 95.99
MFCLIP (Ours) - - 55.96 65.76 76.52 99.93
FSLSD Xception [7] 50.26 54.22 - - 51.49 72.14
ViT [8] 50.01 49.08 - - 50.21 64.98
CViT [10] 50.03 47.60 - - 50.46 84.44
DIRE [9] 50.00 51.13 - - 50.14 57.44
FreqNet [12] 53.44 55.08 - - 49.22 72.31
CLIP [21] 49.67 45.77 - - 52.37 72.90
FatFormer [29] 52.02 47.33 - - 54.71 73.09
VLFFD[15] 51.78 53.56 - - 53.82 84.66
DD-VQA [16] 51.39 54.76 - - 54.60 86.51
MFCLIP (Ours) 53.65 55.59 - - 55.52 92.15
FaceSwapper Xception [7] 49.99 63.86 56.60 45.21 - -
ViT [8] 49.83 42.26 49.72 41.21 - -
CViT [10] 50.00 49.83 50.00 51.97 - -
DIRE [9] 50.00 78.89 50.00 65.41 - -
FreqNet [12] 50.01 46.86 49.82 46.43 - -
CLIP [21] 49.50 44.68 49.01 45.70 - -
FatFormer [29] 51.67 46.83 50.71 47.04 - -
VLFFD[15] 50.00 65.82 51.13 63.48 - -
DD-VQA [16] 50.02 65.96 52.80 59.26 - -
MFCLIP (Ours) 52.03 79.07 59.08 67.23 - -
IV Experiments
To investigate the performance of our MFCLIP detector, we conducted experiments using five FFD datasets: faceforensics++ (FF++) [22], deepfake detection challenge (DFDC) [30], Celeb-DF [31], deeperforensics (DF-1.0) [32], and GenFace [33]. The remainder of this section is organized as follows: 1) the experiment setup including implementation details and datasets is introduced. 2) the performance assessments and comparisons based on five FFD datasets are outlined.
IV-A Experiment Setup
Implementation details. We developed the detector using PyTorch on the Tesla V100 GPU with 32GB memory. The number of blocks $B$ and $L$ in MFCLIP is set to 3 and 6, respectively. The patch size $p$, feature dimension $d$, and batch size $b$ are set to 112, 512, and 24, respectively. We set the channel $c$, height $h$, and width $w$ to 64, 8, and 8, respectively. Our model is trained with the Adam optimizer [34] with a learning rate of 1e-4 and weight decay of 1e-3. We used the scheduler to drop the learning rate by ten times every 15 epochs. We used accuracy (ACC) and area under the receiver operating characteristic curve (AUC) as evaluation metrics.
Datasets. We employed the fine-grained face forgery dataset GenFace [33] for DFFD, where we conducted the cross-forgery protocol and cross-generator evaluation. We used four deepfake datasets to evaluate the generalization and robustness of networks: FF++ [22], DFDC [30], Celeb-DF [31], and DF-1.0 [32].
TABLE III: Cross-diffusion generalization. ACC and AUC scores (%) on remaining diffusion-based generators, after training using one diffusion-based generator.
Training Set Model Testing Set
DDPM LatDiff CollDiff DiffFace Diffae
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
DDPM Xception [7] - - 50.00 63.57 50.48 75.94 53.07 96.80 50.07 87.49
ViT [8] - - 50.67 54.28 52.94 74.53 61.31 88.58 50.21 49.17
CViT [10] - - 50.00 74.51 50.07 73.24 51.51 96.21 50.28 92.62
DIRE [9] - - 50.18 62.50 50.21 69.95 56.54 94.33 51.12 88.55
FreqNet [12] - - 50.00 49.09 50.21 71.52 50.26 91.94 53.77 91.76
CLIP [21] - - 51.79 57.63 52.28 60.95 76.71 93.23 49.86 45.66
FatFormer [29] - - 53.70 59.43 54.66 63.21 78.02 94.80 51.65 48.27
VLFFD [15] - - 51.73 72.96 53.70 74.91 77.46 95.11 52.83 88.87
DD-VQA[16] - - 51.45 73.99 53.84 76.08 77.93 96.05 54.01 89.36
MFCLIP (Ours) - - 55.69 79.80 55.78 77.09 88.89 99.99 55.94 98.76
LatDiff Xception [7] 50.13 76.45 - - 50.00 37.05 51.51 96.21 50.28 92.62
ViT [8] 66.60 80.74 - - 53.20 54.09 52.93 57.59 57.60 65.33
CViT [10] 51.73 90.76 - - 50.00 42.36 50.01 44.50 47.88 44.04
DIRE [9] 50.01 46.29 - - 50.02 43.13 50.01 48.51 50.05 45.98
FreqNet [12] 78.50 83.32 - - 79.04 89.58 50.18 89.93 38.16 37.87
CLIP [21] 62.33 74.53 - - 52.99 53.23 55.71 61.05 56.64 64.38
FatFormer [29] 63.99 76.26 - - 54.90 56.14 57.89 63.04 58.92 66.03
VLFFD [15] 87.35 92.64 - - 56.32 67.81 87.24 89.25 68.53 93.02
DD-VQA[16] 88.46 94.00 - - 57.84 73.02 88.61 97.05 70.56 93.70
MFCLIP (Ours) 99.99 $9$ 65.08 77.07 99.98 99.98 97.92 99.99
CollDiff Xception [7] 55.69 96.31 49.98 70.45 - - 50.17 71.98 50.19 59.50
ViT [8] 55.06 61.83 51.03 48.64 - - 50.26 50.32 50.51 49.97
CViT [10] 99.74 99.97 49.98 47.59 - - 81.97 98.80 90.56 99.83
DIRE [9] 91.52 81.05 50.02 65.99 - - 60.79 93.68 56.85 96.86
FreqNet [12] 93.48 85.42 49.91 55.59 - - 50.04 61.18 99.95 99.98
CLIP [21] 51.58 53.92 50.21 46.17 - - 50.26 48.64 49.82 48.36
FatFormer [29] 53.81 55.52 53.90 49.03 - - 54.80 52.71 52.90 50.65
VLFFD [15] 94.25 96.01 56.89 64.06 - - 82.29 94.67 72.43 92.99
DD-VQA[16] 93.20 97.44 55.60 70.76 - - 83.77 91.59 73.08 92.94
MFCLIP (Ours) 100 100 99.19 99.97 - - 99.63 99.96 93.94 99.99
DiffFace Xception [7] 99.98 99.98 50.00 48.31 50.00 52.97 - - 50.07 87.49
ViT [8] 87.00 97.01 49.11 42.97 49.82 49.00 - - 48.74 40.90
CViT [10] 50.08 77.27 49.98 54.22 50.04 59.65 - - 47.52 40.44
DIRE [9] 50.04 75.32 50.00 47.21 59.79 96.91 - - 73.55 96.53
FreqNet [12] 49.85 82.21 51.77 73.37 49.70 75.62 - - 43.43 58.69
CLIP [21] 76.34 91.74 49.98 50.59 51.56 57.34 - - 55.90 51.90
FatFormer [29] 79.54 94.20 52.89 53.94 53.76 59.02 - - 57.89 54.64
VLFFD [15] 77.84 94.76 62.33 84.97 53.66 60.23 - - 58.08 87.96
DD-VQA[16] 90.22 99.98 63.75 83.88 54.07 59.01 - - 59.89 88.40
MFCLIP (Ours) 99.99 99.99 85.32 99.94 50.57 75.40 - - 52.12 99.92
Diffae Xception [7] 53.96 94.01 49.98 61.27 50.12 68.04 49.99 52.69 - -
ViT [8] 50.52 50.04 49.45 45.06 49.70 47.99 49.51 46.50 - -
CViT [10] 57.78 98.04 50.23 83.37 50.00 54.42 50.13 80.12 - -
DIRE [9] 57.45 94.21 50.07 62.30 50.12 74.84 64.14 99.02 - -
FreqNet [12] 53.48 94.58 49.91 44.41 59.86 53.10 50.04 48.82 - -
CLIP [21] 50.52 51.58 49.91 50.11 50.80 52.07 49.97 50.32 - -
FatFormer [29] 53.56 54.85 52.01 53.89 52.58 54.05 52.57 53.65 - -
VLFFD [15] 62.79 93.43 60.77 82.54 52.78 65.00 63.29 98.36 - -
DD-VQA[16] 64.45 95.06 61.78 83.11 52.87 69.57 60.99 96.47 - -
MFCLIP (Ours) 98.99 99.99 99.82 99.98 60.07 75.80 99.99 $9$
IV-B Comparison with the State of the Art
We evaluate the performance of state-of-the-art deepfake detectors on diffusion-generated images using GenFace. We selected various detectors such as CNN-based Xception [7] and FreqNet [12], transformer-based ViT [8], CViT [10] and FatFormer [29], vision-language-based CLIP [21], VLFFD [15] as well as DD-VQA [16], and DIRE [9] detecting diffusion-generated images.
Figure 7: The robustness of models to unseen various image perturbations.
TABLE IV: Cross-generator evaluation on AM. ACC and AUC scores (%) on remaining generators, after training using one generator.
Training Set Model Testing Set
Diffae LatTrans IAFaces
ACC AUC ACC AUC ACC AUC
Diffae Xception [7] - - 50.00 68.76 52.05 55.51
ViT [8] - - 55.35 76.10 50.45 54.20
CViT [10] - - 50.00 52.03 49.95 66.64
DIRE [9] - - 50.00 41.69 50.30 63.01
FreqNet [12] - - 50.21 51.06 50.00 63.91
CLIP [21] - - 50.00 58.83 49.80 49.65
FatFormer [29] - - 50.34 59.67 50.93 52.45
VLFFD [15] - - 51.02 60.74 51.02 72.81
DD-VQA [16] - - 51.36 69.04 53.84 74.90
MFCLIP (Ours) - - 59.96 79.76 54.55 90.27
LatTrans Xception [7] 51.31 63.14 - - 50.05 50.07
ViT [8] 49.73 49.03 - - 50.05 50.97
CViT [10] 50.76 62.08 - - 50.25 51.71
DIRE [9] 50.02 52.35 - - 50.00 56.72
FreqNet [12] 49.88 49.16 - - 50.00 60.38
CLIP [21] 50.02 47.69 - - 50.00 53.05
FatFormer [29] 50.05 50.21 - - 50.09 54.87
VLFFD [15] 50.07 63.76 - - 50.23 65.96
DD-VQA [16] 52.65 65.00 - - 50.12 66.45
MFCLIP (Ours) 53.00 75.00 - - 55.26 75.41
TABLE V: Cross-dataset generalization. ACC and AUC scores on FF++, Celeb-DF, DFDC, and DF-1.0 after training using FF++.
Method FF++ Celeb-DF DFDC DF-1.0
ACC AUC ACC AUC ACC AUC ACC AUC
ViT[8] 62.44 67.07 62.28 59.75 56.18 58.31 58.05 61.27
CViT[10] 90.47 96.69 50.75 64.70 60.95 65.96 56.15 62.42
RECCE[35] 97.06 99.32 - 68.71 - 69.06 - -
CEViT[36] 93.67 98.36 44.24 65.29 66.14 75.55 62.16 67.51
FoCus[37] 96.43 99.15 - 76.13 - 68.42 - -
UIA-ViT[38] - 99.33 - 82.41 - 75.80 - -
Yu et al.[39] - 99.55 - 72.86 - 69.23 - -
Guan et al.[40] - 99.17 - 95.14 - 74.65 - -
CLIP[21] 67.79 69.57 64.18 65.42 58.42 57.65 57.63 56.01
VLFFD[15] - 99.23 - 84.80 - 84.74 - -
MFCLIP 98.15 99.63 74.02 83.46 79.36 86.08 70.47 78.99
Cross-forgery evaluation. To investigate the generalization of various detectors, we performed cross-forgery tests. We trained models using images of one manipulation, and tested them on those of remaining manipulations. As Table I shows, the performance of our methods outperforms most detectors, demonstrating the superior generalization of MFCLIP. Specifically, for the vision-language-based models, the AUC of our model is nearly 26.80%, 18.78%, and 14.66% higher than that of CLIP, VLFFD, and DD-VQA, respectively, on EFS after training using AM, which is attributed to the powerful learning capabilities of our model for fine-grained text-guided multi-modal face forgery representations.
Cross-generator evaluation. To conduct an in-depth study of the proposed network for generalizable DFFD, we performed the cross-diffusion evaluation on GenFace. Since DFFD is a new and challenging task, to the best of our knowledge, there have been no comprehensive experiments to evaluate the performance of FFD models on cross-diffusion generators. Therefore, our work is the first to assess the generalization of models to unseen diffusion-generated facial images, systematically and comprehensively. Specifically, we trained models using the images generated by one diffusion-based generator and tested them on different ones. As Table III shows, detectors tend to achieve better performance on images produced by the generator with high similarity to the generator used for training. Specifically, since DDPM and DiffFace resemble each other, the detector trained using DDPM shows excellent results on images synthesized by DiffFace, and vice versa. For instance, the AUC of Xception, CViT and DIRE is 96.80%, 96.21%, and 94.33% on DiffFace after training using DDPM, respectively. By contrast, they only acquire 63.57% AUC, 74.51% AUC, and 62.50% AUC on LatDiff, individually.
In addition, as Fig. 6 displays, networks generally perform worse on the images generated by LatDiff than that synthesized by other generators such as DDPM, CollDiff, DiffFace, and Diffae. We believed that facial images produced by LatDiff are realistic, such that the detector struggles to distinguish their authenticity. Therefore, those generated by LatDiff pose a huge threat to the detector. We further performed the cross-generator protocol on AM and FS, respectively. As Table IV and Table II display, most models acquire poor performance (about 60% AUC) on the cross-generator evaluation. The AUC of our network is around 40.60%, 27.26%, and 26.36% higher than that of CLIP, DIRE, and FreqNet, respectively, on IAFaces after training using Diffae, which is attributed to the powerful modelling capability of our MFCLIP method.
Cross-dataset evaluation. To investigate the generalization of MFCLIP, we conducted cross-dataset evaluations. We trained detectors using FF++ and tested them on FF++, CelebDF, DFDC, and DF-1.0. The first two level text prompts in FF++ are only introduced to train MFCLIP, due to the limitation of labels offered by FF++. As Table V shows, the AUC of our MFCLIP method is about 1.34% higher than that of VLFFD, showing the superior generalization ability of MFCLIP.
Robustness to common image corruptions. We assessed the robustness of detectors against different unseen image distortions. We trained models on GenFace and tested their performance on distorted images from [32]. Seven types of perturbations are involved, each with five intensity levels. As shown in Fig. 7, we tested the models on various image distortions, such as saturation changes, contrast adjustments, block distortions, white Gaussian noise, blurring, pixelation, and video compression. We averaged the AUC scores of the detector on seven types of corrupted images. An intensity of 0 indicates no degradation. When adopting perturbations of different severities, the changes in AUC for all detectors are presented in Fig. 7. The results show that our model consistently surpasses all other methods across seven types of image degradation.
Figure 8: (a) and (b) show the ablation results of CLIP and MFCLIP with hierarchical fine-grained text prompts, respectively.
TABLE VI: MFCLIP ablation. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training using CollDiff.
Model Testing Set
DDPM LatDiff DiffFace Diffae
NE IE FTE Predictor SPA ACC AUC ACC AUC ACC AUC ACC AUC
- - - - 99.62 98.83 88.78 88.60 91.06 98.93 89.79 99.86
- - - - 99.74 99.97 49.98 47.59 81.97 98.80 90.56 99.83
- - - 99.77 99.99 92.77 95.48 92.56 99.77 92.35 99.86
- - - 99.76 99.99 92.00 93.14 92.45 99.86 92.11 99.96
- - 99.98 99.99 92.07 95.04 95.48 99.93 92.17 99.97
- 99.99 99.99 93.42 96.18 98.08 99.87 92.18 99.98
100 100 99.19 99.97 99.63 99.96 93.94 99.99
V Ablation Study
We conducted an ablation study to evaluate the contribution of each component in MFCLIP. In this study, we used images generated by five diffusion-based models including CollDiff, DDPM, LatDiff, DiffFace and Diffae, to investigate the effectiveness of MFCLIP. We considered five schemes: 1) the impacts of components, 2) the influence of sample pair attention, 3) the effects of fine-grained text prompts, 4) the influences of loss functions and 5) the effect of the patch size in PS. In the subsequent subsection, we discussed the five aspects, respectively.
V-A Impacts of components
To examine the contribution of each component to learning ability, we observed the performance of models on DDPM, LatDiff, DiffFace, and Diffae after training using CollDiff. Table VI shows the ablation results of the model. NE improves the performance by 47.89% AUC on LatDiff, confirming that noises extracted from the richest patch offer valuable information to benefit the DFFD. The gains from introducing the FLE module (+45.55%) are obvious, demonstrating the significance of fine-grained text embeddings. Predictor increases performance by 3.23% ACC on DiffFace, showing that DFFD could benefit from aligning visual and linguistic representations in feature space. The introduction of SPA further enhances the performance (+3.79%), which verifies that cross-modality sample pairs are adaptively emphasized and suppressed, guiding the model to achieve better feature alignment.
TABLE VII: Ablation results of various patch sizes. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff.
Patch Size DDPM LatDiff DiffFace Diffae
ACC AUC ACC AUC ACC AUC ACC AUC
16 56.21 81.69 50.00 41.57 50.42 68.81 57.31 87.46
28 80.89 99.76 50.00 42.69 50.49 75.34 53.95 87.94
32 98.42 99.96 50.04 49.62 64.49 93.48 63.27 89.95
56 56.30 97.49 50.00 90.79 53.50 97.49 51.59 96.90
112 100 100 99.19 99.97 99.63 99.96 93.94 99.99
224 100 100 95.30 99.83 99.60 99.95 65.50 92.03
V-B Effect of the patch size in PS
We investigated the impact of different patch sizes in PS. The proposed MFCLIP model was trained using CollDiff and tested on DDPM, LatDiff, DiffFace, and Diffae. We reported the performance of MFCLIP from patch size 16 to 224. Table VII shows that the performance typically increases with the growth of patch size. The AUC reaches the peak as the patch size is 112, and begins to decline afterward. We argued that the model tends to explore more forgery areas, when the larger patch size is leveraged. However, oversized patch scales like 224 may introduce noises, enabling the model to acquire poor generalization performance.
TABLE VIII: The effect of the SPA module. We tested models on DDPM, LatDiff, and Diffae, after training using CollDiff.
Method DDPM LatDiff Diffae Params (M) FLOPs (G)
ACC AUC ACC AUC ACC AUC
CLIP w/o SPA 51.58 53.92 50.21 46.17 49.82 48.36 84.225 117.23
CLIP w/ SPA 52.68 59.20 50.37 48.89 49.98 50.21 84.225 117.23
MFCLIP w/o SPA 99.99 99.99 93.42 96.18 92.18 99.98 93.834 358.12
MFCLIP w/ SPA 100 100 99.19 99.97 93.94 99.99 93.834 358.12
Figure 9: (a) The visualization of the cosine similarity matrix in MFCLIP (w/ SPA and w/o SPA). (b) The heatmap visualizations of various detectors on some examples from GenFace.
V-C Influence of sample pair attention
To evaluate the effectiveness and efficiency of SPA, we conducted the cross-generator evaluation to show that our plug-and-play SPA module can generalize to various vison-language-based models for performance improvement. We trained models using CollDiff and tested them on DDPM, LatDiff, and Diffae. As Table VIII shows, due to the addition of MAS, the AUC of both CLIP and MFCLIP is improved by 2.72% and 3.78%, respectively, on the challenging LatDiff. Furthermore, SPA rarely introduces auxiliary weights and computational costs when plugged and played into other models.
V-D Effects of fine-grained text prompts
To investigate the impact of the hierarchical fine-grained text prompts, we evaluated the performance of CLIP and our MFCLIP method on DDPM, LatDiff, and Diffae after training on CollDiff, by gradually introducing the hierarchical texts. In Fig. 8, we noticed that the performance of our model commonly improves with the increase of hierarchical text prompts. Specifically, the AUC of CLIP and our model is improved on DDPM, LatDiff, and Diffae, respectively, as fine-grained texts from level 1 to level 4 are introduced, showing that hierarchical text prompts guide the model to capture more general and discriminative information, to facilitate the advancement of DFFD.
TABLE IX: Effects of losses. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff.
Loss Function DDPM LatDiff DiffFace Diffae
ACC AUC ACC AUC ACC AUC ACC AUC
$\mathcal{L}_{\text{ce}}$ 99.77 99.99 92.77 95.48 92.56 99.77 92.35 99.86
$\mathcal{L}_{\text{ce}}$+$\mathcal{L}_{\text{kl}}$ 99.89 99.99 94.07 97.26 94.53 99.80 92.91 99.92
$\mathcal{L}_{\text{ce}}$+$\mathcal{L}_{\text{cmc}}$ 99.92 99.99 95.36 98.03 95.88 99.85 92.98 99.95
$\mathcal{L}_{\text{ce}}$+$\mathcal{L}_{\text{cmc}}$+$\mathcal{L}_{\text{kl}}$ 100 100 99.19 99.97 99.63 99.96 93.94 99.99
V-E Influences of loss functions
To verify the contribution of loss functions, we performed ablations on various losses. The ablation result of losses is displayed in Table IX. As MFCLIP is guided with merely cross-entropy loss, the AUC is 95.48% on LatDiff, but an about 1.8% increase of AUC could be reached via adding the KL loss. We believed that it could enhance the visual forgery representations via language guidance. Meanwhile, due to the introduction of the CMC loss, MFCLIP is grown by 2.6% AUC on LatDiff, showing the effectiveness of adaptive cross-modal pairs alignment. The integration of the three losses performs the best among these losses, which suggests that the proposed loss could acquire promising results.
VI Visualization
Visualization of SPA. To demonstrate the effectiveness of our SPA method, we visualized the cosine similarity matrix, when SPA is involved or not. As Fig. 9 (a) displays, relevant negative pairs acquire larger correlation scores and vice versa, due to the addition of SPA, which shows that SPA emphasizes the related negative pairs and suppresses the irrelevant ones to achieve flexible alignment.
Visualization of heatmap. To further investigate the effect of MFCLIP, we displayed the heatmap of different detectors in Fig. 9 (b). Each row shows a forgery face yielded by various generators. The second to fifth columns illustrate heatmaps of four models: (I) CLIP; (II) DIRE; (III) FreqNet, and (IV) MFCLIP. Compared to other detectors, MFCLIP (III) captures more manipulated areas, showing that text-guided image-noise face forgery representation learning can benefit FFD. Specifically, in the last row, the pristine face is added with bangs through the attribute-manipulated model Diffae, and we noticed that the MFCLIP model could, to a large extent, identify the forgery area.
VII Conclusion
We propose a novel MFCLIP method to facilitate the advancement of generalizable DFFD. First, we build the fine-grained text generator to produce the text prompts of each image in GenFace. Second, we observe that the significant discrepancy between the authentic and forgery facial SRM noises extracted from the richest patches, compared to the poorest patches, we design the noise encoder to capture the discriminative and fine-grained noise forgery patterns from the richest patches. Furthermore, we devise the fine-grained language encoder to extract the abundant text embeddings. We also present a novel plug-and-play SPA method to align features of cross-modal sample pairs, adaptively, which could be integrated into any vision-language-based model like CLIP with only a slight growth in computational costs.
Limitations. Although our model has explored fine-grained multi-modal forgery traces, we may need to improve generalization across diverse state-of-the-art datasets, and reduce computational complexity. In the future, we intend to integrate the pre-trained MVE with the multi-modal large language model, to improve the ability to understand and detect complex face forgeries.